# CubeDAgger: Improved Robustness of Interactive Imitation Learning without Violation of Dynamic Stability

### v1

@@ Abstract
Interactive imitation learning makes an agent’s control policy robust by stepwise supervisions from an expert. The recent algorithms mostly employ expert-agent switching systems to reduce the expert’s burden by limitedly selecting the supervision timing. However, the precise selection is difficult and such a switching causes abrupt changes in actions, damaging the dynamic stability. This paper therefore proposes a novel method, so-called CubeDAgger, which improves robustness while reducing dynamic stability violations by making three improvements to a baseline method, EnsembleDAgger. The first improvement adds a regularization to explicitly activate the threshold for deciding the supervision timing. The second transforms the expert-agent switching system to an optimal consensus system of multiple action candidates. Third, autoregressive colored noise to the actions is introduced to make the stochastic exploration consistent over time. These improvements are verified by simulations, showing that the learned policies are sufficiently robust while maintaining dynamic stability during interaction.
@@ I Introduction
Robots are expected to automatically perform a wide variety of tasks. The recent paradigm of training large models (i.e. control policies) by collecting a large set of demonstration datasets of tasks [1] can be considered a type of imitation learning [2], so the success of this approach is theoretically promised by minimizing approximation errors as much as possible. However, the costs of collecting data and learning grow as the system grows in size. Collecting enough datasets and acquiring compact policies is still useful in many domains (e.g. those where data is limited [3] and non-stationarity is strong [4]).
Interactive imitation learning (IIL) is a well-known methodology for improving policies by incrementally providing demonstration data [5]. In particular, dataset aggregation (DAgger) [6], which is the focus of this study, allows an agent and an expert to share control of a target robotic system. The expert takes the initiative at the beginning and gradually delegates control to the agent, who can complete the imitation without over- or under-learning. By introducing an appropriate safety criterion for delegating control, the expert can correct the robot to the optimal trajectory before it falls into a high-risk situation [7, 8, 9, 10]. Such a kind of error recovery motions can make the agent policy robust to accumulated approximation errors and/or disturbances.
However, recent DAgger variants have a method of discretely switching control authority, which causes two problems. First, if the expert is a human, he/she may not be able to respond immediately to the switching timing, resulting in a delay in the teaching operation [11]. Alternatively, if the switching timing is handled by him/her [7], the switching timing itself will be delayed. Another problem is that switching controllers is prone to cause discrete and abrupt changes in control commands to the robot, which could easily violate the stability of the robot [12]. Both of these problems can be ignored if the task is limited to (relatively) static manipulations, but in the case of dynamial systems, they easily violate dynamic stability in a synergistic manner.
To resolve this issue, this paper makes three improvements to EnsembleDAgger [8]. First, the agent policy is explicitly regularized so that it satisfies a specified level of safety for determining the control authority. Second, the switching system is replaced with an optimal consensus system of multiple action candidates. Finally, to facilitate more efficient exploration and increase robustness, time-consistent colored noise is injected into the agent’s actions, which is likely to make human decision-making less intractable.
The proposed method, so-called CubeDAgger, with these improvements is tested in three simulation environments with different dynamic characteristics. The results show that the proposed method improves the robustness of the learned policies over the conventional EnsembleDAgger without sacrificing dynamic stability during data collection.
@@ I-A1 DAgger family
DAgger [6] (and its variants [7, 8, 9, 10]) are representative of IIL. An agent and an expert share control of a target robotic system and combine each other’s actions appropriately to safely collect high-quality data. By including the expert’s actions rather than the actual actions into a dataset, the agent policy can asymptotically approach the expert one through supervised learning using it. Alternatively, the expert can improve the agent policy by only directing the agent to modify its actions.
In recent years, it has become common for the way actions are combined to be a switching system. The switching frequency that requires expert teaching is minimized, reducing the burden on the human expert [13, 14]. In this case, a single human operator can teach multiple robots simultaneously, facilitating data collection on a larger scale [15].
However, such an expert-agent switching system causes a delay between the switching of control authority and the correction by the human, and when the human makes the decision to switch control authority as well [7], limiting the target to systems in which the delay is ignorable [11]. In addition, switching systems would violate the stability of dynamic systems because of the discrete and abrupt changes in control commands to the robot [12]. In other words, this recent approach is inappropriate for IIL on dynamic systems.
@@ II-A Behavioral cloning
At first, behavioral cloning (BC) [2] is introduced since it is a main learning method of a agent’s policy even in IIL. Specifically, BC assumes that the target environment (i.e. a robot and surroundings) is Markovian, namely, the optimal policy, $\pi^{\ast}$, can be explained only with the current state observation, $s\in\mathcal{S}\subset\mathbb{R}^{|\mathcal{S}|}$, which is sampled according to the environment’s state transition probability (or its initial state probability) $p_{e}$. $\pi^{\ast}$ is hidden in an expert, but actions sampled from it at each state, $a\in\mathcal{A}\subset\mathbb{R}^{|\mathcal{A}|}$, can be collected. As a result, a dataset $D=\{(s_{n},a_{n})\}^{N}_{n=1}$ with $N$ pairs is prepared in advance.
The agent’s policy $\pi(a\mid s;\theta)$ with $\theta$ the trainable parameters is optimized towards $\pi^{\ast}$ by minimizing the following Kullback-Leibler divergence.
@@
where the expectation operation was approximated by sampling from $D$ generated from $p_{e}\pi^{\ast}$ by Monte Carlo method. The obtained loss function is minimized by, for example, stochastic gradient descent (e.g. AdaTerm [24] employed in this paper), to optimize $\theta$. Note that this study assumes that $\pi$ is a diagonal normal distribution with a learnable scale for continuous action space.
@@ II-B Interactive imitation learning
In BC, $D$ is prepared and given in advance, but as shown in eq. (1), $D$ is intended to approximate the original expectation operation with $p_{e}\pi^{\ast}$. In other words, if $D$ is insufficient, the approximation error will be large and the imitation will fail eventually due to the accumulation of it. Therefore, IIL switches the BC’s problem setting to a new one that appends data into $D$ online, expecting that the policy trained becomes robust enough to the approximation error by collecting sufficient and necessary $D$.
However, to improve the value of expert data, how an agent would behave in the vicinity of an optimal trajectory (e.g. error recovery motions) is important. This reduces the approximation error efficiently and provide robustness to recover from deviations from the optimal trajectory due to the accumulation of errors. The pseudocode of DAgger family [6, 7, 8, 9, 10] with such a function is summarized in Alg. 1.
@@
At Line 7 in Alg. 1, $a^{c}$, which actually acts on the environment, is determined based on $a$ demonstrated by the expert and $a^{\pi}$ of the agent. There is no unique solution for determining $a^{c}$, so several methods have been proposed in previous studies. In the original DAgger [6], the weighted average with annealing was employed. On the other hand, the recent trend is regarded as the following expert-agent switching system.
@@ II-C EnsembleDAgger
As the baseline for the CubeDAgger proposed in the next section, EnsembleDAgger [8] is introduced here. In EnsembleDAgger, the agent evaluate the safety for switching $a$ and $a^{\pi}$, and the expert provides $a$ at evey time step, alleviating the delay of human decision making.
Specifically, EnsembleDAgger has $K$ ensembles of policies, $\{\pi_{k}(a\mid s;\theta_{k})\}_{k=1}^{K}$, each of which is trained with eq. (1). Therefore, the agent has $K$ action candidates, $\{a^{\pi_{k}}\}_{k=1}^{K}$, each of which is given as the location parameter of $\pi_{k}$ (without stochastic exploration) usually. At this time, not only $a^{\pi}=K^{-1}\sum_{k}a^{\pi_{k}}$ but also its standard deviation $\sigma^{\pi}=\sqrt{K^{-1}\sum_{k}(a^{\pi}-a^{\pi_{k}})^{2}}$ can be computed. EnsembleDAgger assumes that the current situation is safe if $a^{\pi}$ is close to $a$ and if $\sigma^{\pi}$ is enough small. With two thresholds, $\bar{\Delta}\geq 0$ and $\bar{\sigma}\geq 0$, the safety set, $\mathcal{C}$, is defined as follows [^1] In the original, both criteria are squared and the thresholds are set accordingly, but this is omitted to make it consistent with the proposal.:
$\mathcal{C}$ =$=\{(s,a)\mid|a-a^{\pi}(s)|\leq\bar{\Delta}\land\sigma^{\pi}(s)\leq\bar{\sigma}\}$ (3)
@@ III-A Overview
    Controlled: To reduce the hand-tuned $\mathcal{C}$ errors, the ensemble uncertainty is explicitly controlled to/under the specified thresholds.
@@ III-B1 Formulation
wherre, $\epsilon\ll 1$ denotes the tiny constant that is mainly used for numerical stabilization. This constraint prevents all $a^{\pi_{k}}$ from perfectly matching the expert’s action $a$, which would impair the statistical performance of the ensemble models, while the upper bound limits the deviation of any action candidate from $a$ to within the threshold. In other words, it is expected to reduce misjudgments of safety while ensuring the effectiveness of the ensemble models.
@@
where, the subscript $i$ refers to the $i$-th dimension of action, and since the inequality constraint is given on each dimension independently, different Lagrange multipliers, $\lambda_{i}$, are provided for each. This implementation simplifies learning without the use of RL algorithms as in the literature [25].
@@ III-B2 Optimization for satisfying inequality constraint
$\mathcal{L}^{\delta}(\phi_{\delta})$ =$=\begin{cases}\delta(s_{n};\phi_{\delta})\mathrm{sign}(e(a,a^{\pi},\delta))&|e|>\bar{e}\\ \delta(s_{n};\phi_{\delta})\lambda&\mathrm{otherwise}\end{cases}$ (8)
where, $g_{\delta}(\cdot)\in[0,\ln(1+\epsilon)-\ln(2\epsilon)]$ is the nonlinear transformation to satisfy the domain of the slack variable. $\bar{e}$ denotes the allowable error, and in this paper, it is designed as 10 % of the domain size. If $e\leq\bar{e}$, the position of equality constraint is shifted towards the lower or upper bound of $\delta$ according to $\lambda$: if $\lambda>0$ to maximize the ensemble uncertainty, its behavior is suppressed by making $\delta$ small, and vice versa. This behavior allows the change in the ensemble uncertainty to be stably controlled while simultaneously searching for solutions that may lie at both boundaries of the inequality constraint.
@@ III-C Consensus system among ensemble actions
Recent DAgger variant, such as EnsembleDAgger, use a safety-based switching system to decide which action $a^{c}$ to execute from the expert’s action candidate $a$ and the agent’s action candidate $a^{\pi}$. However, as mentioned above, safety decisions left to human are subject to delay, and switching can cause abrupt changes in robot’s commands, which can violate the stability of the dynamic system. Therefore, this study considers the decision of $a^{c}$ as a kind of consensus-making problem in which a central tendency is derived from multiple candidates, and designs a system that can obtain an optimal consensus with safety taken into account.
@@ III-C1 Formulation
where, $\underline{a}_{i}$ and $\overline{a}_{i}$ denote the minimum and maximum values among the candidates, respectively. $w_{k}$ is the weigtht for $k$-th candidate, the sum of which satisfies $1$. Note that when $a^{c}$ is inferred only by agents (especially after deployment), the $K+1$-th candidate for $a$ is merely excluded.
@@
The above properties are the same even when $w_{k}$ is included, since it only vary the number of pseudo-candidates. Thus, given appropriate weights, the conventional discrete switching system can be extended to a continuous one.
@@ III-C2 Design of $p_{i}$
This ratio is known to be $\sqrt{2/\pi}(\simeq 0.8)$ if the candidates follow a Gaussian distribution. If the candidates follow a uniform distribution, it asymptotically approaches $1$, and if they are non-Gaussian, such as heavily tailed and/or asymmetric, it approaches 00. Therefore, $p_{i}$ can be defined as the function of $\rho_{i}$. However, since $p_{i}$ and $\rho_{i}$ have different domains, a reasonable connection between them is required.
This paper focuses on the longest distance $d_{i}$ from the origin among sets in a $K+1$-dimensional space where the $L_{p_{i}}$ norm is $1$ (see Fig. 1(a)). Three important cases are as follows: $\sqrt{K+1}^{-1}$ for $p_{i}\to 1$; $1=\sqrt{K+1}^{0}$ for $p_{i}=2$; and $\sqrt{K+1}^{1}$ for $p_{i}\to\infty$. Here, the power part (defined as $q_{i}$) is bounded by $(-1,1)$, and the connection with $\rho_{i}$ can be easily made by the following equation.
$d_{i}=\sqrt{K+1}^{q_{i}}=2\rho_{i}^{\frac{\ln 2}{\ln\sqrt{\pi/2}}}-1$ (11)
@@ III-D Colored noise for time-consistent exploration
In the context of RL, action noise from the stochastic policy is effective, and DART variants [16, 17], another approach to IIL, make the learned policy robust by adding optimized noise to the expert’s behavior. Thus, the exploration noise is also considered effective even in DAgger variants. In the above research cases (with continuous action space), however, the exploration noise is white noise generated from a Gaussian distribution. In other words, noise is time-independent, and a robot system fastly oscillated by it might be difficult for a human expert to operate.
@@
$a^{\pi}=\mu_{\pi}+\cfrac{2\sqrt{K}}{3}\sigma_{\pi}\epsilon^{\mathrm{red}}_{t}$ (15)
where, $\mu_{\pi}$ and $\sigma_{\pi}$ denote the mean and scale of $\pi$, which is modeled as a diagonal normal distribution, respectively. Note that $\epsilon^{\mathrm{red}}_{t}$ is different for each dimension of the action space and for each component of the ensemble model. As white and red noises are generally within ±$\pm 2\sim 3$, this gain, $32\sqrt{K}/$ 3, increases the expert’s confidence level relatively only to $K\sim 2K$, even if all components carry the maximum amount of exploration noise. In other words, the difference of the number of candidates from the agent and expert is compensated with this design, and the actions from the agent are possibly reflected into $a^{c}$ (depending on $p_{i}$ and $w_{k}$).
IV Simulations

### v2

# CubeDAgger: Interactive Imitation Learning for Dynamic Systems with Efficient yet Low-risk Interaction

@@ Abstract
Interactive imitation learning makes an agent’s control policy robust by stepwise supervisions from an expert. The recent algorithms mostly employ expert-agent switching systems to reduce the expert’s burden by limitedly selecting the supervision timing. However, this approach is useful only for static tasks; in dynamic tasks, timing discrepancies cause abrupt changes in actions, losing the robot’s dynamic stability. This paper therefore proposes a novel method, named CubeDAgger, which improves robustness with less dynamic stability violations even for dynamic tasks. The proposed method is designed on a baseline, EnsembleDAgger, with three improvements. The first adds a regularization to explicitly activate the threshold for deciding the supervision timing. The second transforms the expert-agent switching system to an optimal consensus system of multiple action candidates. Third, autoregressive colored noise is injected to the agent’s actions for time-consistent exploration. These improvements are verified by simulations, showing that the trained policies are sufficiently robust while maintaining dynamic stability during interaction. Finally, real-robot scooping experiments with a human expert demonstrate that the proposed method can learn robust policies from scratch based on just 30 minutes of interaction.
@@ I Introduction
Interactive imitation learning (IIL) is a well-known methodology for improving policies by incrementally providing demonstration data [6]. In particular, dataset aggregation (DAgger) [28], which is the focus of this study, allows an agent and an expert to share control of a target robotic system. Conceptually, the expert takes the initiative at the beginning and gradually delegates control to the agent, who can complete the imitation without over- or under-learning. By introducing an appropriate safety criterion for delegating control, the expert can correct the robot to the optimal trajectory before it falls into a high-risk situation [16, 22, 8, 23]. Such an error recovery motion can make the agent policy robust to accumulated approximation errors and/or disturbances.
Recent DAgger variants commonly have a method of discretely switching control authority, which causes two problems. First, if the expert is a human, he/she may not be able to respond immediately to the switching timing, causing a delay in the teaching operation [3]. Even if the switching timing is handled by him/her [16], the switching timing itself would be delayed. These delays are not particularly problematic for static tasks (that is why IIL mostly targets a static object manipulation), but they can be fatal for dynamic tasks. For example, errouneous switching controllers cause discrete and abrupt changes in control commands to the robot, which could easily violate the robot’s dynamic stability [20].
To enable IIL for dynamic tasks, this paper proposes a new Dagger variant, named CubeDAgger. The proposed CubeDAgger (or C3DAgger) has three ‘C’ improvements on EnsembleDAgger [22]. First, the agent policy is explicitly Controlled so that it satisfies a specified level of safety for determining the control authority. Second, the switching system is replaced with an optimal Consensus system of multiple action candidates. Finally, time-consistent Colored noise is injected into the agent’s actions for efficient exploration, which also makes human decision-making less intractable.
The proposed CubeDAgger is tested in three simulation environments with different dynamic characteristics. The results show that CubeDAgger improves the robustness of the trained policies over the conventional EnsembleDAgger without sacrificing dynamic stability during data collection. Additionally, a real-world scooping task (see Fig. 1) is demonstrated. Although baselines make human supervision errouneous and imitation unsatisfactory, CubeDAgger can acquire robust policies in just 30 minutes of interaction while maintaining comfortable supervision.
Figure 1: Scooping three balls by a net attached on a quadruped
@@ I-A1 DAgger family
DAgger [28] (and its variants [16, 22, 8, 23]) are representative of IIL. An agent and an expert share control of a target robotic system and combine each other’s actions appropriately to safely collect high-quality data. By including the expert’s actions rather than the actual actions into a dataset, the agent policy can asymptotically approach the expert one through supervised learning using it.
In recent years, a switching system gained popularity as a way to combine actions of the agent and the expert. The switching frequency that requires expert teaching is minimized, reducing the burden on the human expert [12, 11]. In this case, a single human operator can teach multiple robots simultaneously, facilitating data collection on a larger scale [13].
However, such an expert-agent switching system causes a delay between the switching of control authority and the correction by the human, and when the human makes the decision to switch control authority as well [16], limiting the target to systems in which the delay is ignorable [3]. In addition, switching systems would violate the stability of dynamic systems because of the discrete and abrupt changes in control commands to the robot [20]. In other words, this recent approach is unsuitable for IIL on dynamic systems.
@@ II-A Behavioral cloning
At first, behavioral cloning (BC) [2] is introduced since it is a main learning method of an agent’s policy even in IIL. Specifically, BC assumes that the target environment (i.e. a robot and surroundings) is Markovian, namely, the optimal policy, $\pi^{\ast}$, can be expressed only with the current state observation, $s\in\mathcal{S}\subset\mathbb{R}^{|\mathcal{S}|}$, which is sampled according to the environment’s state transition probability $p_{e}$ (or its initial state probability $p_{0}$). $\pi^{\ast}$ is hidden in an expert, but actions sampled from it at each state, $a\in\mathcal{A}\subset\mathbb{R}^{|\mathcal{A}|}$, can be collected. As a result, a dataset $D=\{(s_{n},a_{n})\}^{N}_{n=1}$ with $N$ pairs is prepared in advance.
The agent’s policy $\pi(a\mid s;\theta)$ with $\theta$ the trainable parameters is optimized towards $\pi^{\ast}$ by minimizing the following loss function.
@@
This problem is solved by, for example, stochastic gradient descent (e.g. AdaTerm [14] employed in this paper). Note that this study assumes that $\pi$ is a diagonal normal distribution with a learnable scale for continuous action space.
@@ II-B Interactive imitation learning
In BC, $D$ is prepared and given in advance, but $D$ is intended to approximate the original expectation operation with $p_{e},\pi^{\ast}$. In other words, if $D$ is insufficient, the approximation error will be large and the imitation will fail eventually due to the accumulation of it. Therefore, IIL switches the BC’s problem to a new one that appends data into $D$ online, making the policy trained robust enough to the approximation error by collecting sufficient and necessary $D$.
In addition, to improve the value of expert data, how an agent would behave in the vicinity of an optimal trajectory (e.g. error recovery motions) is important. This reduces the approximation error efficiently and provide robustness to recover from deviations from the optimal trajectory due to the accumulation of errors. The pseudocode of DAgger family [28, 16, 22, 8, 23] with such a function is summarized in Alg. 1.
@@
At Line 7 in Alg. 1, $a^{c}$, which actually acts on the environment, is determined based on $a$ demonstrated by the expert and $a^{\pi}$ of the agent. There is no unique solution for determining $a^{c}$, so several methods have been proposed in previous studies. In the original DAgger [28], the weighted average with annealing was employed. On the other hand, the recent trend is the following expert-agent switching system.
@@ II-C EnsembleDAgger
As the baseline for the CubeDAgger proposed in the next section, EnsembleDAgger [22] is introduced here. In EnsembleDAgger, the agent evaluates the safety for switching $a$ and $a^{\pi}$, and the expert provides $a$ at evey time step, alleviating the delay of human decision making.
Specifically, EnsembleDAgger has $K$ ensembles of policies, $\{\pi_{k}(a\mid s;\theta_{k})\}_{k=1}^{K}$, each of which is trained with eq. (1). Therefore, the agent has $K$ action candidates, $\{a^{\pi_{k}}\}_{k=1}^{K}$, each of which is given as the location parameter of $\pi_{k}$ (without stochastic exploration). At this time, not only $a^{\pi}=K^{-1}\sum_{k}a^{\pi_{k}}$ but also its standard deviation $\sigma^{\pi}=\sqrt{K^{-1}\sum_{k}(a^{\pi}-a^{\pi_{k}})^{2}}$ can be computed. EnsembleDAgger assumes that the current situation is safe if $a^{\pi}$ is close to $a$ and if $\sigma^{\pi}$ is enough small. With two thresholds, $\bar{\Delta}\geq 0$ and $\bar{\sigma}\geq 0$, the safety set, $\mathcal{C}$, is defined as follows [^1] In the original, both criteria are squared and the thresholds are set accordingly, but this is omitted to make it consistent with the proposal.:
$\mathcal{C}$ $=\{(s,a)\mid|a-a^{\pi}(s)|\leq\bar{\Delta}\land\sigma^{\pi}(s)\leq\bar{\sigma}\}$ (3)
@@ III-A Overview
    Controlled: To reduce the errors of the hand-tuned safety set, the ensemble uncertainty is explicitly controlled to/under the specified thresholds.
@@ III-B1 Formulation
where, $\epsilon\ll 1$ denotes the tiny constant that is mainly used for numerical stabilization. This constraint prevents all $a^{\pi_{k}}$ from perfectly matching the expert’s action $a$, which would impair the statistical performance of the ensemble models, while the upper bound limits the deviation of any action candidate from $a$ to within the threshold.
@@
where, the subscript $i$ refers to the $i$-th dimension of action, and since the inequality constraint is given on each dimension independently, different Lagrange multipliers, $\lambda_{i}$, are provided for each. If the policies are trained well, the inquality constraint should be satisfied correctly. In other words, it is expected to reduce misjudgments of safety while ensuring the effectiveness of the ensemble models.
@@ III-B2 Optimization for satisfying inequality constraint
$\mathcal{L}^{\delta}(\phi_{\delta})$ $=\begin{cases}\delta(s_{n};\phi_{\delta})\mathrm{sign}(e(a,a^{\pi},\delta))&|e|>\bar{e}\\ \delta(s_{n};\phi_{\delta})\lambda&\mathrm{otherwise}\end{cases}$ (8)
where, $g_{\delta}(\cdot)\in[0,\ln(1+\epsilon)-\ln(2\epsilon)]$ is the nonlinear transformation to satisfy the domain of the slack variable. $\bar{e}$ denotes the allowable error, and in this paper, it is designed as 10 % of the domain size, referring to the literature [18]. If $e\leq\bar{e}$, the position of equality constraint is shifted towards the lower or upper bound of $\delta$ according to $\lambda$: if $\lambda>0$ to maximize the ensemble uncertainty, its behavior is suppressed by making $\delta$ small, and vice versa. This behavior allows the change in the ensemble uncertainty to be stably controlled while simultaneously searching for solutions that may lie at both boundaries of the inequality constraint.
@@ III-C Consensus system among ensemble actions
Recent DAgger variants, such as EnsembleDAgger, use a safety-based switching system to decide which action $a^{c}$ to execute from the expert’s action candidate $a$ and the agent’s action candidate $a^{\pi}$. However, as mentioned above, safety decisions left to human are subject to delay, and switching can cause abrupt changes in robot’s commands, which can violate the stability of dynamic systems. Therefore, this study considers the decision of $a^{c}$ as a kind of consensus-making problem in which a central tendency is derived from multiple candidates, and designs a system that can obtain an optimal consensus with safety taken into account.
@@ III-C1 Formulation
where, $\underline{a}_{i}$ and $\overline{a}_{i}$ denote the minimum and maximum values among the candidates, respectively. $w_{k}\geq 0$ is the weigtht for $k$-th candidate, the sum of which satisfies $1$. Note that when $a^{c}$ is inferred only by agents (especially after deployment), the $K+1$-th candidate for $a$ is merely excluded.
@@
The above properties are the same even when $w_{k}$ is included, since it only vary the pseudo-number of candidates. Thus, given appropriate weights, the conventional discrete switching system can be extended to a continuous one.
@@ III-C2 Design of $p_{i}$
This ratio is known to be $\sqrt{2/\pi}(\simeq 0.8)$ if the candidates follow a Gaussian distribution. If the candidates follow a uniform distribution, it asymptotically approaches $1$, and if they are non-Gaussian, such as heavily tailed and/or asymmetric, it approaches 0. Therefore, $p_{i}$ can be defined as the function of $\rho_{i}$. However, since $p_{i}$ and $\rho_{i}$ have different domains, a reasonable connection between them is required.
This paper focuses on the longest distance $d_{i}$ from the origin among sets in a $K+1$-dimensional space where the $L_{p_{i}}$ norm is $1$ (see Fig. 2(a)). Three important cases are as follows: $\sqrt{K+1}^{-1}$ for $p_{i}\to 1$; $1=\sqrt{K+1}^{0}$ for $p_{i}=2$; and $\sqrt{K+1}^{1}$ for $p_{i}\to\infty$. Here, the power part (defined as $q_{i}$) is bounded by (1,1)(-1,1), and the connection with $\rho_{i}$ can be easily made by the following equation.
$q_{i}=2\rho_{i}^{\frac{\ln 2}{\ln\sqrt{\pi/2}}}-1$ (11)
@@ III-D Colored noise for time-consistent exploration
In the context of RL, action noise from the stochastic policy is effective, and DART variants [19, 24], another approach to IIL, make the trained policy robust by adding optimized noise to the expert’s behavior. Thus, the exploration noise is also considered effective even in DAgger variants. In the above research cases (with continuous action space), however, the exploration noise is white noise generated from a Gaussian distribution. In other words, noise is time-independent, and a robot system fastly oscillated by it might be difficult for a human expert to operate (see the attached video and the section IV-B).
@@
$a^{\pi_{k}}=\mu_{\pi_{k}}+\cfrac{2\sqrt{K}}{3}\sigma_{\pi_{k}}\epsilon^{\mathrm{red}}_{t,k}$ (15)
where, $\mu_{\pi_{k}}$ and $\sigma_{\pi_{k}}$ denote the mean and scale of $\pi_{k}$, which is modeled as a diagonal normal distribution, respectively. Note that the component-wise $\epsilon^{\mathrm{red}}_{t,k}$ is different for each dimension of the action space. As white and red noises are generally within $\pm 2\sim 3$, this gain, $2\sqrt{K}/3$, increases the expert’s confidence level relatively only to $K\sim 2K$, even if all components carry the maximum amount of exploration noise. In other words, the difference of the number of candidates from the agent and expert is compensated with this design, and the actions from the agent are possibly reflected into $a^{c}$ (depending on $p_{i}$ and $w_{k}$).
IV Experiments